# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v2

@@ Abstract.
We study the dynamic query evaluation problem: Given a natural join query $Q$ and a stream of updates, we would like to construct a data structure that supports constant-delay enumeration of the query output after each update.
We show that a stream of $N$ insert-only updates to an initially empty database can be executed in total time $\mathcal{O}(N^{\mathsf{w}(Q)})$, where $\mathsf{w}(Q)$ is the fractional hypertree width of $Q$. This matches the complexity of the static query evaluation problem for $Q$ and a database of size $N$. One corollary is that the amortized time per single-tuple insert is constant for acyclic joins.
@@ 1. Introduction
Answering queries under updates to the underlying database, also called dynamic query evaluation, is a fundamental problem in data management, with recent work addressing it from both systems and theoretical perspectives. There have been recent efforts on building systems for dynamic query evaluation, such as DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017, 2020), and F-IVM (Nikolic and Olteanu, 2018; Nikolic et al., 2020; Kara et al., 2023b). By allowing tuples to carry payloads with elements from rings (Koch, 2010), such systems can maintain complex analytics over database queries, such as linear algebra computation (Nikolic et al., 2014), collection programming (Koch et al., 2016), and machine learning models (Kara et al., 2023b). Beyond conjunctive queries, there has been work on intersection joins under updates (Tao and Yi, 2022), and dynamic computation for expressive languages such as Datalog (Motik et al., 2019), Differential Datalog (Abadi et al., 2015), or DBSP (Budiu et al., 2023). The descriptive complexity of various recursive Datalog queries under updates, such as reachability, has also been investigated (Datta et al., 2015, 2018; Schwentick et al., 2020).
@@
Third, we devise new dynamic query evaluation algorithms that use our reductions. They improve or recover best known update times, e.g., amortized constant update time for any acyclic query in the insert-only setting and for any hierarchical query in the insert-delete setting. They also achieve optimal $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized update time for Loomis-Whitney queries in the insert-delete setting, where $\mathcal{D}$ is the database size at update time.
@@ 1.1. Motivating Examples
[…]
We show that $Q_{p}$ admits single-tuple updates: in amortized constant time in the insert-only setting; and in amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ time in the insert-delete setting ($\tilde{O}$ notation hides a $\text{\sf polylog}(|\mathcal{D}|)$ factor). Both times are optimal, the latter is conditioned on the OMv conjecture. The amortized constant time per insert does not contradict the non-constant lower bound for non-hierarchical queries without self-joins (Berkholz et al., 2017): that lower bound proof relies on a reduction from OMv that requires both inserts and deletes. Furthermore, it is non-trivial to achieve amortized constant time per insert, while also allowing constant-delay enumeration after each insert; such a result was only shown recently in (Wang et al., 2023). The lazy maintenance approach, which only updates the input data, is not sufficient to achieve this result. This is because, while it does indeed take constant insert time, it still requires the computation of the query output before the enumeration of the first output tuple. This computation takes linear time using a factorized approach (Kara et al., 2023b) and needs to be repeated for every enumeration request.
The triangle query $Q_{\triangle}$ and the Loomis-Whitney-4 query $Q_{4}$ need $\mathcal{O}(|\mathcal{D}|^{1/2})$ and $\mathcal{O}(|\mathcal{D}|^{1/3})$, respectively, amortized single-tuple insert times. The total times of $\mathcal{O}(|\mathcal{D}|^{3/2})$ and $\mathcal{O}(|\mathcal{D}|^{4/3})$, respectively, match the worst-case optimal times for the static evaluation of these two queries (Ngo et al., 2018). In the insert-delete setting, single-tuple updates take $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized time for both queries, and this is optimal based on the OMv conjecture.
@@ 2. Problem Setting
Problem: $\text{IVM}^{\pm}[Q]$ Parameter: Query $Q$ Given: An initially empty database $\mathcal{D}$ and a stream $\delta_{1}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$. $N$ is not known in advance. Task: Support constant-delay enumeration of $Q(\mathcal{D}^{(\tau)})$ after each update $\delta_{\tau}\mathcal{D}$
@@
Problem: $\text{IVM}^{\pm}_{\delta}[Q]$ Parameter: Query $Q$ Given: An initially empty database $\mathcal{D}$ and a stream $\delta_{1}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$. $N$ is not known in advance. Task: Support constant-delay enumeration of the change $\delta_{\tau}Q(\mathcal{D})$ after each update $\delta_{\tau}\mathcal{D}$
@@ 3 $+T(a_{1},c_{1})$$\{(a_{1},b_{1},c_{1})\}$$\{+Q(a_{1},b_{1},c_{1})\}$
Table 1. A stream of 8 updates to $Q_{\triangle}$ in Eq. (1). The last two columns show the full and delta result after each update.
@@
Problem: $\text{Eval}[Q]$ Parameter: Query $Q$ Given: Database $\mathcal{D}$ Task: Compute $Q(\mathcal{D})$ For this problem, we are interested in the time to compute $Q(\mathcal{D})$.
@@
The function $\mathsf{\omega}(Q)$ is not known in general. However, we know that $\mathsf{\omega}(Q)\leq\mathsf{w}(Q)$; see e.g. (Abo Khamis et al., 2016). More strongly, $\mathsf{\omega}(Q)$ is upper bounded by the submodular width of $Q$ (Marx, 2013; Abo Khamis et al., 2017)222 (Abo Khamis et al., 2017) provides a query evaluation algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)}\cdot\text{\sf polylog}{|\mathcal{D}|}+|Q(\mathcal{D})|)=\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)+o(1)}+|Q(\mathcal{D})|)$, thanks to the extra $+o(1)$ in the exponent. $\mathrm{subw}(Q)$ denotes the submodular width of $Q$. . For acyclic queries, we have $\mathsf{\omega}(Q)=1$ (Yannakakis, 1981). For the triangle query, $\mathsf{\omega}(Q_{\triangle})\geq\frac{4}{3}$ modulo the 3SUM conjecture (Patrascu, 2010). In this paper, we prove several lower bounds on the IVM problems in terms of the function $\mathsf{\omega}(Q)$. These lower bounds are not conditioned on any fine-grained complexity conjecture (like the OMv conjecture (Henzinger et al., 2015), stated in Appendix A). Instead, they apply no matter what the value of $\mathsf{\omega}(Q)$ turns out to be. In that sense, they can be thought of as unconditional.
@@
For all problems introduced above, we consider the query to be fixed. Moreover, we say that the amortized update time for the $\tau$-th single-tuple update is $f(N,|\mathcal{D}^{(\tau)}|)$ for some function $f$, if the total time to process all $N$ updates in the stream is upper bounded by $\sum_{\tau\in[N]}f(N,|\mathcal{D}^{(\tau)}|)$. In the insert-only setting, the database size $|\mathcal{D}^{(\tau)}|$ at time $\tau$ is always equal to $\tau$. Hence, for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$, we ignore the database size and only measure the update time as a function of the length $N$ of the update stream. In the insert-delete setting, the database size at time $\tau$ is upper bounded by $\tau$ and could be much smaller than $\tau$. As a result, for $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$, it is more natural to ignore $N$ and only measure the update time as a function of $|\mathcal{D}^{(\tau)}|$. For the four IVM problems, the constant enumeration delay is not allowed to be amortized, regardless of whether the update time is amortized or not. For the problem Eval, we measure the computation time as a function of the size of the input database $\mathcal{D}$. We use classical $\mathcal{O}$-notation to express worst-case bounds and $\widetilde{\mathcal{O}}$-notation to hide a polylogarithmic factor in $N$, $|\mathcal{D}^{(\tau)}|$, or respectively $|\mathcal{D}|$.
@@
For the purpose of stating our results, we introduce the following concept.
@@
$\widehat{Q}_{\bm{\sigma}}=\widehat{R}_{\sigma_{1}}(Z_{1},\bm{X}_{\sigma_{1}})$ $\wedge\widehat{R}_{\sigma_{2}}(Z_{1},Z_{2},\bm{X}_{\sigma_{2}})$
(2) $\wedge\cdots\wedge\widehat{R}_{\sigma_{k}}(Z_{1},\ldots,Z_{k},\bm{X}_{\sigma_{k}})$
@@
(4) $\hskip-11.38092pt\widehat{Q}_{123}=\widehat{R}_{123}(Z_{1},A,B)\wedge\widehat{S}_{123}(Z_{1},Z_{2},B,C)\wedge\widehat{T}_{123}(Z_{1},Z_{2},Z_{3},A,C)$
@@ 3.1. Insert-Only Setting
In the insert-only setting, we show that off-the-shelf worst-case optimal join algorithms, e.g., LeapFrog TrieJoin (Veldhuizen, 2014), can be used to achieve the best known amortized update time for the dynamic query evaluation of arbitrary join queries, so including the cyclic queries, while supporting constant delay for full/delta enumeration. We give an upper bound on the single-tuple update time in terms of the fractional hypertree width $\mathsf{w}(Q)$ of $Q$:
Theorem 3.1 ().
[…]
Sec. 5 overviews our algorithm that meets the upper bound in Theorem 3.1. It involves worst-case optimal join algorithms and tree decompositions. Amortization is necessary as some inserts may be costly, while many others will necessarily be relatively cheaper, so the average insert cost matches the cost of computing a factorized representation of the query result divided by the number of inserts.
@@
[…]
Corollary 3.2 recovers the amortized constant update time for acyclic joins from prior work (Wang et al., 2023). It also shows that the insert-only setting can be computationally cheaper than the insert-delete setting investigated in prior work (Berkholz et al., 2017): In the insert-delete setting, the non-hierarchical acyclic queries cannot admit $\mathcal{O}(|\mathcal{D}|^{1/2-\gamma})$ update time for any database $\mathcal{D}$ and $\gamma>0$ (Berkholz et al., 2017), conditioned on the Online Matrix-Vector Multiplication (OMv) conjecture.
We accompany the above upper bound with a lower bound on the insertion time in terms of the lower bound $\mathsf{\omega}(Q)$ on the static evaluation of $Q$. This lower bound is not conditioned on any fine-grained complexity conjecture (like OMv). Instead, it holds no matter what the value of $\omega(Q)$ is:
Theorem 3.3 ().
[…]

### v3

@@ Abstract.
We study the dynamic query evaluation problem: Given a join query $Q$ and a sequence of updates, we would like to construct a data structure that supports constant-delay enumeration of the query output after each update.
We show that a sequence of $N$ insert-only updates to an initially empty database can be executed in total time $\mathcal{O}(N^{\mathsf{w}(Q)})$, where $\mathsf{w}(Q)$ is the fractional hypertree width of $Q$. This matches the complexity of the static query evaluation problem for $Q$ and a database of size $N$. One corollary is that the amortized time per single-tuple insert is constant for acyclic joins.
@@ 1. Introduction
Answering queries under updates to the underlying database, also called dynamic query evaluation, is a fundamental problem in data management, with recent work addressing it from both systems and theoretical perspectives. There have been recent efforts on building systems for dynamic query evaluation, such as DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017, 2020), F-IVM (Nikolic and Olteanu, 2018; Nikolic et al., 2020; Kara et al., 2023b), and CROWN (Wang et al., 2023). By allowing tuples to carry payloads with elements from rings (Koch, 2010), such systems can maintain complex analytics over database queries, such as linear algebra computation (Nikolic et al., 2014), collection programming (Koch et al., 2016), and machine learning models (Kara et al., 2023b). Beyond conjunctive queries, there has been work on intersection joins under updates (Tao and Yi, 2022), and dynamic computation for expressive languages such as Datalog (Motik et al., 2019), Differential Datalog (Abadi et al., 2015), or DBSP (Budiu et al., 2023). The descriptive complexity of various recursive Datalog queries under updates, such as reachability, has also been investigated (Datta et al., 2015, 2018; Schwentick et al., 2020).
@@
Third, we devise new dynamic query evaluation algorithms that use our reductions. We call them MVIVM (short for MultiVariate IVM). They improve or recover best known update times, e.g., amortized constant update time for any acyclic query in the insert-only setting and for any hierarchical query in the insert-delete setting. They also achieve optimal $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized update time for Loomis-Whitney queries in the insert-delete setting, where $\mathcal{D}$ is the database size at update time.
@@ 1.1. Motivating Examples
We would like to perform single-tuple updates to each query, while ensuring constant-delay enumeration of the tuples in the query output after each update. Theses queries are not $q$-hierarchical, hence they cannot admit worst-case constant time per single-tuple update (Berkholz et al., 2017): The OMv conjecture implies that there is no algorithm that takes $\mathcal{O}(|\mathcal{D}|^{1/2-\gamma})$ time for single-tuple updates, for any $\gamma>0$, to any of these queries. In fact, for a database of size $\mathcal{D}$, systems like DBToaster (Koch et al., 2014), DynYannakakis (Idris et al., 2017), CROWN (Wang et al., 2023), and F-IVM (Nikolic and Olteanu, 2018), need at least $\mathcal{O}(|\mathcal{D}|)$ time for a single-tuple update to each of these queries. For the triangle query $Q_{\triangle}$, IVMϵ takes $\mathcal{O}(|\mathcal{D}|^{1/2})$ amortized time for single-tuple updates (Kara et al., 2020a). In the insert-only setting, CROWN maintains $Q_{3p}$ in amortized constant time per insert. However, it cannot handle non-acyclic queries like $Q_{\triangle}$ and $Q_{4}$ (with update time better than recomputation from scratch). Appendix A explains how each prior approach maintains $Q_{3p}$ and $Q_{\triangle}$ in the insert-delete setting, and compares their update times to our approach, MVIVM.
We show that, for $Q_{3p}$, MVIVM performs single-tuple updates: in amortized constant time in the insert-only setting; and in amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ time in the insert-delete setting ($\tilde{O}$ notation hides a $\text{\sf polylog}(|\mathcal{D}|)$ factor). Both times are optimal, the latter is conditioned on the OMv conjecture. The amortized constant time per insert matches that of CROWN (Wang et al., 2023) and does not contradict the non-constant lower bound for non-hierarchical queries without self-joins (Berkholz et al., 2017): that lower bound proof relies on a reduction from OMv that requires both inserts and deletes. Achieving this amortized constant time, while also allowing constant-delay enumeration after each insert, is non-trivial. The lazy maintenance approach, which only updates the input data, is not sufficient to achieve this result. This is because, while it does indeed take constant insert time, it still requires the computation of the query output before the enumeration of the first output tuple. This computation takes linear time using a factorized approach (Kara et al., 2023b) and needs to be repeated for every enumeration request.
MVIVM takes $\mathcal{O}(|\mathcal{D}|^{1/2})$ and $\mathcal{O}(|\mathcal{D}|^{1/3})$ amortized single-tuple insert time for the triangle query $Q_{\triangle}$ and, respectively, the Loomis-Whitney-4 query $Q_{4}$. Given a sequence of $|\mathcal{D}|$ inserts, the total times of $\mathcal{O}(|\mathcal{D}|^{3/2})$ and $\mathcal{O}(|\mathcal{D}|^{4/3})$, respectively, match the worst-case optimal times for the static evaluation of these two queries (Ngo et al., 2018). In the insert-delete setting, single-tuple updates take $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized time for both queries, and this is optimal based on the OMv conjecture. Figure 3 in Appendix A shows the update times of MVIVM and prior work for further queries.
@@ 2. Problem Setting
Problem: $\text{IVM}^{\pm}[Q]$ Parameter: Query $Q$ Given: An initially empty database $\mathcal{D}$ and a stream $\delta_{1}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$ where $N$ is not known in advance. Task: Support constant-delay enumeration of $Q(\mathcal{D}^{(\tau)})$ after each update $\delta_{\tau}\mathcal{D}$
@@
Problem: $\text{IVM}^{\pm}_{\delta}[Q]$ Parameter: Query $Q$ Given: An initially empty database $\mathcal{D}$ and a stream $\delta_{1}\mathcal{D},\ldots,\delta_{N}\mathcal{D}$ of $N$ single-tuple updates to $\mathcal{D}$ where $N$ is not known in advance. Task: Support constant-delay enumeration of $\delta_{\tau}Q(\mathcal{D})$ after each update $\delta_{\tau}\mathcal{D}$
@@ 3 $+T(a_{1},c_{1})$$\{(a_{1},b_{1},c_{1})\}$$\{+Q(a_{1},b_{1},c_{1})\}$
Table 1. An update sequence for the query $Q_{\triangle}$ in Eq. (1). The last two columns show the full and delta result after each update.
@@
Problem: $\text{Eval}[Q]$ Parameter: Query $Q$ Given: Database $\mathcal{D}$ Task: Compute $Q(\mathcal{D})$
For this problem, we are interested in the time to compute $Q(\mathcal{D})$.
@@
The function $\mathsf{\omega}(Q)$ is not known in general. However, we know that $\mathsf{\omega}(Q)\leq\mathsf{w}(Q)$; see e.g. (Abo Khamis et al., 2016). More strongly, $\mathsf{\omega}(Q)$ is upper bounded by the submodular width of $Q$ (Marx, 2013; Abo Khamis et al., 2017; Abo Khamis et al., 2024)222 (Abo Khamis et al., 2017; Abo Khamis et al., 2024) provide a query evaluation algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)}\cdot\text{\sf polylog}{|\mathcal{D}|}+|Q(\mathcal{D})|)=\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)+o(1)}+|Q(\mathcal{D})|)$, thanks to the extra $+o(1)$ in the exponent. $\mathrm{subw}(Q)$ denotes the submodular width of $Q$. . For acyclic queries, we have $\mathsf{\omega}(Q)=1$ (Yannakakis, 1981). For the triangle query, $\mathsf{\omega}(Q_{\triangle})\geq\frac{4}{3}$ modulo the 3SUM conjecture (Patrascu, 2010). In this paper, we prove several lower bounds on the IVM problems in terms of the function $\mathsf{\omega}(Q)$. These lower bounds are not conditioned on any fine-grained complexity conjecture (like the OMv conjecture (Henzinger et al., 2015), stated in Appendix B). Instead, they apply no matter what the value of $\mathsf{\omega}(Q)$ turns out to be. In that sense, they can be thought of as unconditional. [^3] Alternatively, these lower bounds can be thought of as conditioned on (symbolic) lower bounds on the static query evaluation problem.
@@
For all problems introduced above, we consider the query to be fixed. Moreover, we say that the amortized update time for the $\tau$-th single-tuple update is $f(N,|\mathcal{D}^{(\tau)}|)$ for some function $f$, if the total time to process all $N$ updates in the stream is upper bounded by $\sum_{\tau\in[N]}f(N,|\mathcal{D}^{(\tau)}|)$. In the insert-only setting, the database size $|\mathcal{D}^{(\tau)}|$ at time $\tau$ is always equal to $\tau$. Hence, for $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$, we ignore the database size and only measure the update time as a function of the length $N$ of the update stream. In the insert-delete setting, the database size at time $\tau$ is upper bounded by $\tau$ and could be much smaller than $\tau$. As a result, for $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$, it is more natural to ignore $N$ and only measure the update time as a function of $|\mathcal{D}^{(\tau)}|$. For the four IVM problems, the constant enumeration delay is not allowed to be amortized, regardless of whether the update time is amortized or not. For the problem Eval, we measure the computation time as a function of the size of the input database $\mathcal{D}$. We use the $\mathcal{O}$-notation to state worst-case bounds and $\widetilde{\mathcal{O}}$-notation to hide a polylogarithmic factor in $N$, $|\mathcal{D}^{(\tau)}|$, or $|\mathcal{D}|$.
@@
For the purpose of stating our results, we introduce the following central concept, which gives our approach its name, MultiVariate IVM (MVIVM for short).
@@
(2) $\widehat{Q}_{\bm{\sigma}}=\widehat{R}_{\sigma_{1}}(Z_{1},\bm{X}_{\sigma_{1}})\wedge\widehat{R}_{\sigma_{2}}(Z_{1},Z_{2},\bm{X}_{\sigma_{2}})\wedge\cdots\wedge\widehat{R}_{\sigma_{k}}(Z_{1},\ldots,Z_{k},\bm{X}_{\sigma_{k}})$
@@
(4) $\widehat{Q}_{123}=\widehat{R}_{123}(Z_{1},A,B)\wedge\widehat{S}_{123}(Z_{1},Z_{2},B,C)\wedge\widehat{T}_{123}(Z_{1},Z_{2},Z_{3},A,C)$
@@ 3.1. Insert-Only Setting
In the insert-only setting, we show that off-the-shelf worst-case optimal join algorithms, e.g., LeapFrog TrieJoin (Veldhuizen, 2014), can be used to achieve the best known amortized update time for the dynamic query evaluation of arbitrary join queries, so including the cyclic queries, while supporting constant delay for full/delta enumeration. Specifically, we give an upper bound on the single-tuple update time in terms of the fractional hypertree width $\mathsf{w}(Q)$ of $Q$:
Theorem 3.1.
For any query $Q$, both $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\mathcal{O}(N^{\mathsf{w}(Q)-1})$ amortized update time, where $N$ is the number of single-tuple inserts.
The upper bound in Theorem 3.1 is met by our algorithm outlined in Sec. 5. It uses worst-case optimal join algorithms and tree decompositions. Amortization is necessary as some inserts may be costly, while many others are necessarily relatively cheaper, so the average insert time matches the cost of computing a factorized representation of the query result divided by the number of inserts.
@@
For any acyclic query $Q$, $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\mathcal{O}(1)$ amortized update time.
Corollary 3.2 recovers the amortized constant update time for acyclic joins from prior work (Wang et al., 2023). It also shows that the insert-only setting can be computationally cheaper than the insert-delete setting investigated in prior work (Berkholz et al., 2017): In the insert-delete setting, the non-hierarchical acyclic queries cannot admit $\mathcal{O}(|\mathcal{D}|^{1/2-\gamma})$ update time (while keeping the enumeration delay constant) for any database $\mathcal{D}$ and $\gamma>0$ (Berkholz et al., 2017), conditioned on the OMv conjecture.
We accompany the above upper bound with a lower bound on the insertion time in terms of the lower bound $\mathsf{\omega}(Q)$ on the static evaluation of $Q$. This lower bound is not conditioned on a fine-grained complexity conjecture (like OMv). It holds no matter what the value of $\omega(Q)$ is:
Theorem 3.3.
For any query $Q$ and any constant $\gamma>0$, neither $\text{IVM}^{+}[Q]$ nor $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{\omega}(Q)-1-\gamma})$ amortized update time.